Perform the inverse Laplace transform of the following function. Please state any pairs or properties used

5 85 10 + V, 6) = 10 (1-4*** 5 5² (s +4545.45)

Vo(s) =10^4(1-e^(-0.00000001s))*(85/s^2 + 10^5/s^2*(s+4545.45)
Answer #1

Vors) = 100 (1-e -(0X10%) Learn los 32/5 +4545.45) r] let okox10 9 - A 4546.45 =B + V; («) = 104 (1-0-As) les 105 52/5 +B) )-BT [ é et la é - ) -B -B - Bt ) rlleng] Sot е at 3 = -1 B -35 e -B +). -it I t w 10 1 (nr 1 le -1) t B S 1 left-) - 2 Il Po